OpenStudy (anonymous):

helpp A cylinder has its height doubled and its radius cut to one third. What is the ratio of the volumes of the modified cylinder to the original cylinder?

OpenStudy (wolf1728):

doubling the height doubles the volume if the RADIUS is reduced to 1/3 its size, the volume is reduced by 1/9 so the volume is changed by 2/1 * 1/9 equals 2/9 of the original.

OpenStudy (anonymous):

is there a formula ?

OpenStudy (anonymous):

@DDCamp can you please help me?

OpenStudy (anonymous):

The volume of a cylinder is: \[V_{1} =h \pi r^{2}\] If you double the height (h ---> 2h) You get \[V_{2} = 2h \pi r^{2}\] Which is the double of V1. If you cut the radius to a third (r ----> r/3 ) You get \[V_{3} = 2h \pi \left( \frac{ r }{ 3 } \right)^{2} = \frac{ 2 }{ 9 } h \pi r^{2}\]

OpenStudy (anonymous):

So the Volume is transformed to 2/9ths the original volume (V1).

OpenStudy (anonymous):

Thanks Lessis :)

OpenStudy (anonymous):

No problem. ;D